OpenStudy (anonymous):

Calculate the volume of a cylinder that has a diameter of 8 inches and a height of 3 inches. A. 24 in.³ B. 151 in.³ C. 604 in.³ D. 75 in.³ E. 43 in.³

OpenStudy (anonymous):

Do you know the formula for finding the volume of a cylinder?

OpenStudy (anonymous):

No i don't

OpenStudy (anonymous):

volume of cylinder = (base area)(height) Base area in this question is (8/2)^2 pi Height is 3 so the volume = ((8/2)^2 pi)(3)

OpenStudy (anonymous):

how do you find the base area?

OpenStudy (anonymous):

area of circle = (radius)^2 pi

OpenStudy (anonymous):

the base of a cylinder is a circle in order to the find the area of a circle we use this equation a= pi * r^2 where a is the area and r is the radius In this case you were given the diameter 2*r = d where r represents the radius and d represents the diameter

OpenStudy (anonymous):

I can't figure out the answer I did what you guys said but it didn't work? :(

OpenStudy (phi):

first what did you get for the base area? You can estimate (in your head) like this: \[ \pi r^2 \] is (very roughly) 3* 4*4= 48 (pi is really 3.14159...) the height is 3, so (using 50 instead of 48 'cause it's easier) 3*50= 150 I'd expect the answer about 150 or so

OpenStudy (anonymous):

ok one of the answers is 151 will that work?

OpenStudy (phi):

You should try to do the calculation. Get a calculator. type in 4*4 (that is the radius squared) multiply it by 3.14 that will be the area multiply by the height (3) that should be the volume

OpenStudy (anonymous):

ok thanks